OpenStudy (anonymous):

Find the factored form of -11x-x^2=0

OpenStudy (id21):

\[-x(11+x)=0\]

OpenStudy (id21):

x=0 or x=-11

OpenStudy (anonymous):

-11x-x^2=0 -(11x+x^2)=0 x(11+x)=0 x=0,x=-11